we have 5 gallons of a 8% bleach solution. We want 5 gallons of a 6% bleach solution. How much 8% solution and how much pure water do we need? Thanks!
sueaft@gmail.com

Found 2 solutions by stanbon, jorel1380:
Answer by stanbon(75887)

we have 5 gallons of a 8% bleach solution. We want 5 gallons of a 6% bleach solution. How much 8% solution and how much pure water do we need?
------
Equation:
active - active + active = active
0.08*5 - 0.08x + 0x = 0.06*5
---
Multiply thru by 100 to get:
8*5 - 8x = 6*5
8x = 2*5
---
x = 5/4 gallon (amt 8% solution to be replaced by pure water)
================
Cheers,
Stan H.

Answer by jorel1380(3719)